# DHGS: Decoupled Hybrid Gaussian Splatting for Driving Scene

### v1

@@ Abstract
Existing Gaussian splatting methods struggle to achieve satisfactory novel view synthesis in driving scenes due to the lack of crafty design and geometric constraints of related elements. This paper introduces a novel method called Decoupled Hybrid Gaussian Splatting (DHGS), which aims at promoting the rendering quality of novel view synthesis for driving scenes. The novelty of this work lies in the decoupled and hybrid pixel-level blender for road and non-road layers, without conventional unified differentiable rendering logic for the entire scene, meanwhile maintaining consistent and continuous superimposition through the proposed depth-ordered rendering strategy. Beyond that, an implicit road representation comprised of Signed Distance Field (SDF) is trained to supervise the road surface with subtle geometric attributes. Accompanied by the use of auxiliary transmittance loss and consistency loss, novel images with imperceptible boundary and elevated fidelity are ultimately obtained. Substantial experiments on Waymo dataset prove that DHGS outperforms the state-of-the-art methods.
Figure 1: Comparative analysis of rendered images on novel view synthesis of road models is presented, the top row displaying results without SDF regularizer, and the bottom row showcasing results obtained using SDF regularizer. By visualizing corresponding ellipsoids, it can be observed that the inclusion of SDF regularization leads the road model to learn higher-quality geometric information.
@@ Introduction
The prevalence of 3DGS (Kerbl et al. 2023) has drawn numerous attentions in the community of 3D reconstruction and autonomous driving. Compared with NeRF-based approaches (Mildenhall et al. 2020; Barron et al. 2021), researchers who adopt Gaussian splatting are able to train and deploy a model in a more efficient and flexible way. Although Gaussian splatting technology has achieved immense performance in scene reconstruction, there are still imperfections in the novel view synthesis capabilities, which are particularly crucial for data synthesis tasks of autonomous driving scenes. Existing methods either model the whole driving scene in a consolidated manner (Kerbl et al. 2023; Huang et al. 2024a; Cheng et al. 2024; Lu et al. 2024), resulting in an even growing and pruning strategy for all types of scene elements, i.e. road, building, and distant views, or model the near scene and far sky (Wu et al. 2023; Miao et al. 2024) dividually. Above mentioned approaches emphasize the overall or specific far elements without the attention to nearby synthesis quality. For real-world autonomous driving scenarios, the autonomous vehicle is usually on a smooth road equipped with several cameras with minimal overlapping fields of view, driving at a high speed. We reckon that the road should be maintained as the priority, since road-related geometrics play the basic role in downstream perceptual tasks and are brittle in view changes.
@@
In this paper, we propose a method named Decoupled Hybrid Gaussian Splatting for Driving Scene (DHGS) for lifting the rendering quality of novel view synthesis, especially the nearby surroundings with fewer flaws and more subtle details. The insight of the method comes from the existing geometric prior that could be utilized to supervise the nearby road elements along with other non-road environments in a fancy pattern. Specifically, we decouple the whole driving scene into a nearby road model and environment model represented by two separate Gaussian models, considering optimizing the two models based on their geometric characters accordingly. Before that, we project and segment the point clouds using semantic 2D masks to obtain the initial road point cloud and environmental point cloud. The road point cloud is used for pre-training the implicit Signed Distance Field (SDF). The training process is conducted via the introduced depth-ordered rendering strategy by the blender of rendered images of both models, assisted with transmittance loss that regularizes the accumulated transmittance of each component Gaussian properties, as well as consistency loss which bonds the two areas compactly. Besides, we optimize the road elements using SDF loss that is expressly designed for surface regularization, reaching consistent and orderly road Gaussians. To summarize, the contributions of the proposed DHGS are listed as follows:
@@
    We first propose to decouple the driving scene into near road model and other non-road model, enabling the optimization separately of two models and rendering jointly with the designed depth-ordered rendering strategy,
@@
    We propose an implicit road represent for better guidance of road Gaussian, alone with transmittance loss and consistency loss that guarantee continuity and consistency,
@@
    Abundant quantitative and qualitative experiments on the popular Waymo dataset demonstrate that DHGS achieves state-of-the-art rendering qualities on training views and brand-new free-view synthesis.
@@
NeRF (Mildenhall et al. 2020) has attracted remarkable attention in the field of view synthesis, which adopts Multi-Layer Perceptron (MLP) to implicitly model the 3D scene. NeRF takes the spatial coordinates and the viewing directions of the sample point as input to MLP and predicts the color and opacity of each point in the camera ray, then accumulates into pixel color by neural volume rendering. For each pixel, necessitating the sampling of hundreds of points and subsequent prediction of their attributes via an MLP, leading to extremely slow training and rendering phases. Besides, the finite number of sampling points limits the expression ability in complex and large-scale driving scenes. Mip-NeRF (Barron et al. 2021) proposes a more efficient sampling strategy, where the sampling region is improved from a single ray to a frustum of a viewing cone. Mip-NeRF 360 (Barron et al. 2022) and NeRF++ (Zhang et al. 2020) divide the scene into near and far regions, and apply warping operations to the far regions to handle the unbounded 3D scene. Instant-NGP (Müller et al. 2022) uses a hash grid to store features and reduces the MLP size to release the computational burden. Plenoxels Fridovich-Keil et al. (2022) makes use of sparse grids to reduce storage space and utilizes Spherical Harmonics (SH) to represent the appearance. These methods are based on the data captured by 360-degree object-centric trajectory, while free and long driving scenes draw less attention. F2-NeRF (Wang et al. 2023) proposes a novel space-warping method that supports arbitrary input camera trajectories, enabling the reconstruction of driving scenes. StreetSurf (Guo et al. 2023) extends prior object-centric neural surface reconstruction techniques to address the unique challenges posed by the unbounded street views, showing potential for various downstream tasks. The aforementioned approaches remain a trade-off between rendering quality and efficiency, thereby limiting the application in large-scale autonomous driving scenarios.
@@
3DGS (Kerbl et al. 2023) is a seminal work that pioneeres the use of 3D Gaussian ellipsoids to explicitly represent a scene and leverages CUDA for parallel rendering, achieving rendering quality and speed that far exceed NeRFs. It integrates pixel colors using $\alpha$-blending and performs adaptive density control, thus the input only requires sparse point clouds. A number of works have been proposed for large-scale scene 3D reconstruction based on 3DGS. VastGaussian (Lin et al. 2024) tackles large scenes with novel designs for scene partitioning, optimizing, and merging. HUGS (Zhou et al. 2024a) offers the ability to render novel views in real-time, yielding 2D and 3D semantic information with high accuracy.
@@
Figure 2: The pipline of the proposed method for driving scene reconstruction. For consecutive multi-camera images and corresponding road and non-road masks, we firstly generate decoupled road pcd (point cloud) and environment pcd, a road SDF is then pre-trained as the following supervision for road gaussian model. Environment pcd enables the initialization for environment Gaussian model which composes rendered images with images by paratactic road model via proposed Depth-ordered rendering.
@@
The vanilla 3DGS only fits the consistency of pixel colors, neglecting to model the geometric structure. Consequently, obvious discontinuities such as artifacts and holes appear under unseen views. This discontinuity phenomenon would be more prominent on training datasets with sparse views. DRGS (Chung, Oh, and Lee 2024) introduces a dense depth map as a geometry guide to mitigate overfitting. AtomGS (Liu et al. 2024) introduces an atomized proliferation strategy aimed at enhancing rendering quality by refining 3D geometric precision in areas with fine details. NeuSG (Chen, Li, and Lee 2023) flattens the Gaussian ellipsoid by designing a regularization loss for the scale parameter, and performs consistent joint optimization in combination with the signed distance field. SuGaR (Guédon and Lepetit 2024) first proposes adding a regularization term to encourage the alignment of the Gaussian with the surface, thereby achieving accurate and fast mesh extraction. Scaffold GS (Lu et al. 2024) precisely controls the density of the Gaussian ellipsoid by establishing voxel anchors and corresponding offset vectors, and learning the various parameters of the Gaussian through a simple multi-layer network. Gaussian Pro (Cheng et al. 2024) considers the planar prior in the scene, explicitly constraining the growth of Gaussians, achieving amendatory rendering and more compact representation. 2DGS (Huang et al. 2024a) presents a highly efficient differentiable 2D Gaussian renderer, enabling perspective-accurate splatting by leveraging 2D surface modeling, demonstrating superiority in geometric reconstruction.
Reconstructing a reliable scene geometry plays a pivotal role in driving data synthesis techniques for 3DGS-like approaches. Driving Gaussians (Zhou et al. 2024b), Street Gaussians (Yan et al. 2024) and S3 Gaussians (Huang et al. 2024b) all initialize the point cloud collected by LiDAR instead of the point cloud generated by SfM. With the use of dense and precise point clouds from LiDAR, additional supervision, such as the positional constraint, is introduced to the learning process.
Existing works pay less attention to the ability of novel view synthesis, leading to low-grade image quality when imposed remarkable camera transformation. We propose to address this shortcoming by decoupling the entire scene into two distinct models, employing optimization separately.
@@
where $\mu$ is the mean position of Gaussian and the $x$ is a discretional coordinate within the Gaussian space. $\Sigma=RSS^{T}R^{T}$, where $R$ is the rotation matrix and $S$ is the scaling matrix. Apart from the spherical harmonic coefficients that represent the anisotropic color of each Gaussian, each pixel color is aggregated by $\alpha$-blending:
$\mathbf{c}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
@@
The proposed method utilizes initial points clouds and semantic masks as paratactic input of multi-camera views. PCD initialization first generates road and non-road point clouds that will be further modeled as road and environment Gaussian model. Based on the known road point cloud, we propose to design a mathematical implicit road representation by SDF which serves as a prior for the surface training. The SDF-based surface constraint consists of pre-training and off-line supervision phases, by taking advantage of the distance and normal characters geometrically. We choose to model road and non-road elements with two kinds of Gaussians for better rendering quality in perspective shifts. To achieve that, depth-ordered hybrid rendering is delicately designed, by which the road surface and non-road areas can be coupled and superimposed consistently and continuously, yielding superior performance compared with SOTA methods that employ individual Gaussian model. The rendered image by fusing two models will be supervised by ground truth via Gaussian loss, and optimized in collaboration with the regularization terms. We will go through the process in turn in the following sections.
PCD Initialization
@@
Different from previous approaches, we opt to utilize LiDAR scan which exhibits better multi-view consistency and prior geometric structure compared to the point cloud from SfM. The initial point cloud is classified into road and non-road parts for the subsequent training. As shown in Figure 2, based on the calibrated intrinsic and extrinsic parameters, multi-camera images and their masks generated by Mask2Former (Cheng et al. 2022) are utilized to form colored and semantically labeled single-frame point clouds. Then we stitch these separate point clouds to build road and environment point clouds.
Pre-trained Surface Base On SDF
@@
To optimize the geometric structure of roads and maintain the continuity when facing obvious changes in perspective, we design a road constraint based on surface guidance. Taking advantage of the separated road point cloud, we pre-train a signed distance field as an implicit representation of the real road surface. This manner differs from the joint ray sampling optimization in NeRF and the guidance of geometric structure and image consistency in GSDF (Yu et al. 2024). As road point clouds collected by LiDAR have been calibrated precisely in advance, the excellent geometric structure is suitable to be prior for road learning. A straightforward strategy is directly applying distance constraints to Gaussians that belong to road. Whereas, considering the spatial geometric characteristics of ellipsoids, imposing constraints only on the centers of Gaussian ellipsoids cannot fully restore their correct geometric shapes and may ultimately limit the Gaussian distribution to some extent. We therefore adopt a pre-trained model to optimize the Gaussian distribution by adjusting Gaussian parameters nearing the road surface. To this end, we design several regularization terms specific to road-related Gaussians, including distance constraints and normal direction constraints, forcing Gaussians to be close to the road surface and directional aligned with the road normal synchronously.
@@
We utilize a predefined network $f_{\theta}(x)$ for predicting the corresponding SDF value given the positional coordinate of the point cloud. Data normalization is imposed on all the training and testing coordinates for stable and better performance. For balancing the number of points on the isosurface and off the isosurface, we make training data as follows. Suppose that $\mathcal{P}_{r}$ represents the road point cloud, each point $x_{p}\in\mathcal{P}_{r}$ meanwhile on isosurface will be assigned a hypothetical SDF value of 0. For other points off the isosurface, we set a nearby range for each point in $\mathcal{P}_{r}$ and conduct random sampling within that region, thus obtaining sample points $\mathcal{P}_{s}\supset\mathcal{P}_{r}$, for $\forall x_{s}\in\mathcal{P}_{s}$ we build a map from point $x_{s}$ to the signed euclidean distance $d_{s}$ as:
@@
where $x_{r}^{s}$ is the nearest point of $x_{s}$ in $\mathcal{P}_{r}$. Considering that bumps exist in the uneven road surfaces, we generate the normal direction for each point in $\mathcal{P}_{r}$ to assist in the training the SDF. To be specific, for $\forall x_{r}\in\mathcal{P}_{r}$, set $X_{r}$ as $k$-nearest neighbors of $x_{r}$, by performing SVD:
@@
in which the first and last terms refer to SDF value regularization and eikonal regularizer respectively, the middle term represents normal loss, ||$\cdot$|| ⋅ | refers to the number of points in the point cloud. The optimal parameters $\theta^{*}=\arg\min_{\theta}{L(\theta)}$ will be frozen and used to guide the learning of road Gaussians in the following processes.
@@
Table 1: The quantitative comparison between our method with four SOTA works on PSNR, SSIM, LPIPS and FID metrics computed over all clips on Waymo dataset. We adopt the same training data for each method and follow the corresponding experimental settings strictly. The best and second best are marked in bold and underline for each metric of specific setting, respectively.
Figure 3: The schematic of the proposed depth-ordered hybrid rendering strategy for the environment and road model. Corresponding primitives of each model generate pixels with colors independently through Gaussian splatting. These colors are then composited based on their rendered depths and transmittances, producing the final rendered image.
@@
where $\lambda_{r}$ and $\lambda_{e}$ are the weights for fusing $I_{r}$ and $I_{e}$ rendered from environment model and road model respectivelly, which are organized as:
@@
Figure 4: Comparison of different methods on Waymo dataset, the left column and right column showcase the quality of scene reconstruction and novel view synthesis respectively. Our method achieves high-quality reconstruction for both the environment region and the road region, outperforming other comparative methods in both aspects.
@@
where ${M}$ refers to the semantic mask of the image, ||$\cdot$|| ⋅ | denotes the total number of pixels for given matrix. The $\mathcal{L}_{c}$ enforces consistency between the depth rendered by enviroment gaussians and road gaussians:
@@
where $i$ and $j$ satisfying $\tilde{M}^{i,j}\not=0$, $\tilde{M}$ refers to the banded boundary generated by binary mask $M$, and the orange band at the bottom right corner of Figure 2 intuitively shows the distribution of elements in $\tilde{M}$. We follow Plenoxels (Fridovich-Keil et al. 2022) and use a total variation loss $\mathcal{L}_{\text{tv}}$. Through the joint efforts of abovementioed loss functions, DHGS is able to obtain high-quality render views with continuous geometric structure.
@@
Datasets and Metrics. We experiment on the publicly available urban scene datasets Waymo (Sun et al. 2020), 32 static scenes are adopted as the same as EmerNeRF (Yang et al. 2023). We employ multiple metrics to evaluate the performance of the proposed method and baselines. PSNR, SSIM, and LPIPS metrics are utilized to validate the quality of rendered images with ground truth. For novel views without ground truth, we introduce FID (Heusel et al. 2017) to evaluate the image quality. The averaged metrics across all clips are calculated as comparative values.
Baselines. We compare our method with state-of-the-art and publicly available works. 3DGS (Kerbl et al. 2023), 2DGS (Huang et al. 2024a), Gaussian Pro (Cheng et al. 2024), and Scaffold GS (Lu et al. 2024) are chosen as the comparison methods having regarded to their outstanding performances.
dRefer to caption
Figure 5: Visual comparisons on the free-view novel view synthesis. The left and right columns respectively showcase the results of Set3 and Set4 under the free viewpoint setting, where our method significantly outperforms other comparative methods in both road and environment details.
Implementation details. The initial point cloud is randomly downsampled to 600,000 points. All methods are trained for 60,000 iterations. For the pre-trained SDF network, we employ an 8-layer MLP and train it for 20,000 iterations. The normals of surface points are computed using SVD with 50 neighbors, and the weights for the normal loss and the eikonal loss of the SDF network are set to 0.1 and 0.01. The weights $\lambda_{\text{tran}}$, $\lambda_{\text{sdf}}$, $\lambda_{\text{cons}}$, $\lambda_{\text{tv}}$, $\lambda_{d}$ and $\lambda_{n}$ of the overall training loss are set to 0.1, 1.0, 0.04, 0.1, 0.1 and 0.1. We set the densification intervals for the Gaussian ellipsoids in the environment model and road model to 200 and 300 respectively. The learning rate for anchor position in the environment model is set as 0.00016, the remaining learning rates are kept consistent with the original settings. For each clip, we utilize all the front, left-front, and right-front views of the entire sequence for training, with a 4:1 split ratio between the training and testing sets. The training set is referred to as the reconstruction views, while the testing set is called the novel views. We further apply unrestricted transformations on camera poses to validate the potential of model on data-driven simulation, which is referred to as free-view novel view synthesis.
@@
We compare against baseline methods on scene reconstruction, novel view synthesis, and free-view novel view synthesis that represents notable viewpoints transfer. All methods use the same set of reconstruction views to train and use the rest as novel views to test for each clip. The qualitative and quantitative results will be analyzed in the following sections, where the ablation study is also conducted.

### v2

@@ Abstract
Existing Gaussian splatting methods often fall short in achieving satisfactory novel view synthesis in driving scenes, primarily due to the absence of crafty design and geometric constraints for the involved elements. This paper introduces a novel neural rendering method termed Decoupled Hybrid Gaussian Splatting (DHGS), targeting at promoting the rendering quality of novel view synthesis for static driving scenes. The novelty of this work lies in the decoupled and hybrid pixel-level blender for road and non-road layers, without the conventional unified differentiable rendering logic for the entire scene, while still maintaining consistent and continuous superimposition through the proposed depth-ordered hybrid rendering strategy. Additionally, an implicit road representation comprised of a Signed Distance Field (SDF) is trained to supervise the road surface with subtle geometric attributes. Accompanied by the use of auxiliary transmittance loss and consistency loss, novel images with imperceptible boundary and elevated fidelity are ultimately obtained. Substantial experiments on the Waymo dataset prove that DHGS outperforms the state-of-the-art methods. The project page where more video evidences are given is: https://ironbrotherstyle.github.io/dhgs_web.
Figure 1: The comparison of rendered road images and ellipsoids on novel view synthesis is presented, with the top row displaying results without SDF regularizer, and the bottom row showcasing results obtained using SDF regularizer. It can be observed that the inclusion of SDF regularization leads the road model to render higher-quality images with the help of enhanced road geometry.
@@ Introduction
Autonomous Driving (AD) is a data-driven task that appeals the attention of 3D reconstruction community for ages. Due to the expense of collecting a tremendous amount of data, the neural rendering technique which mainly consists of Neural Radiance Fields (NeRFs) (Mildenhall et al. 2020; Barron et al. 2021) and Gaussian Splatting (Kerbl et al. 2023) emerges as an alternative. With the gathered AD data, a synthetic driving scenario can be generated, facilitating close-loop simulation and testing, with a variety of sensor configurations not typically achievable in real-world tests Yang et al. (2023b); Tonderski et al. (2024); Lindström et al. (2024). Compared with NeRF-based approaches, researchers utilizing Gaussian splatting can train and deploy a model more efficiently and flexibly. Despite the significant advancements in scene reconstruction offered by Gaussian splatting, there are still imperfections in the novel view synthesis capabilities, especially for critical downstream perceptual tasks such as online mapping (Li et al. 2024) of AD scenes. Current methods either model the whole driving scene in a consolidated manner (Kerbl et al. 2023; Huang et al. 2024a; Cheng et al. 2024; Lu et al. 2024), necessitating a uniform growth and pruning strategy for all scene elements, i.e. road, building, and distant views, or model the near scene and far sky (Wu et al. 2023; Miao et al. 2024) dividually. These approaches tend to emphasize the overall or specific distant elements, often overlooking the nearby synthesis quality, which is brittle and significantly impacted by the change of camera perspective. We believe that prioritizing the road is essential, as its geometric properties are fundamental to the success of AD systems.
@@
In this paper, we propose a static neural rendering method named Decoupled Hybrid Gaussian Splatting for Driving Scene (DHGS) to lift the rendering quality of novel view synthesis, particularly for nearby surroundings with fewer flaws and more subtle details. The insight of the method comes from the existing geometric prior that could be utilized to supervise the nearby road elements along with other non-road environments in a fancy pattern. Specifically, we decouple the whole driving scene into a nearby road model and environment model represented by two distinct Gaussian models, considering optimizing the two models based on their geometric characters accordingly. Before that, point clouds are projected and segmented using semantic 2D masks to derive the initial road and environmental point clouds. The road point cloud is used for pre-training the implicit Signed Distance Field (SDF), which serves as the supervision of road training. The training process is conducted via the introduced depth-ordered hybrid rendering strategy by the blender of rendered images of both models, assisted with transmittance loss that regularizes the accumulated transmittance of each componential Gaussian property, as well as consistency loss which bonds the two areas compactly. Besides, we refine the road elements using an SDF loss that is expressly designed for surface regularization, achieving uniform and orderly road Gaussians. To summarize, the contributions of the proposed DHGS are listed as follows:
@@
    We first propose to decouple the driving scene into near road model and other non-road model, enabling the separate optimization of two models and rendering jointly with the designed depth-ordered hybrid rendering.
@@
    We propose an implicit road representation for enhanced guidance of road Gaussians, complemented by transmittance and consistency losses ensuring continuity and consistency.
@@
    Extensive quantitative and qualitative experiments on the popular Waymo dataset demonstrate that DHGS achieves state-of-the-art performance on scene reconstruction and diversified novel free-view synthesis.
@@
NeRF (Mildenhall et al. 2020) has garnered remarkable attention within the neural rendering domain, pioneering the use of a Multi-Layer Perceptron (MLP) to model 3D scenes implicitly. The computationally intensive neural volume rendering and the finite number of sampling points hinder the deployment of NeRF in intricate and expansive driving scenarios. Mip-NeRF (Barron et al. 2021) proposes a more efficient sampling strategy, refining the sampling region from a single ray to a frustum of a viewing cone. Both Mip-NeRF 360 (Barron et al. 2022) and NeRF++ (Zhang et al. 2020) segment the scene into proximal and distal zones, employing warping operations on the latter to handle the unbounded 3D scene. Instant-NGP (Müller et al. 2022) leverages a hash grid for feature storage and streamlines the MLP size to release the computational burden. Plenoxels (Fridovich-Keil et al. 2022) employs sparse grids to reduce storage space and utilize Spherical Harmonics (SH) to represent the appearance. These approaches rely on data from 360-degree object-centric trajectories, while free and long driving scenes draw less attention. F2-NeRF (Wang et al. 2023) proposes a novel space-warping method that supports arbitrary input camera trajectories, enabling the reconstruction of driving scenes. StreetSurf (Guo et al. 2023) extends prior object-centric neural surface reconstruction techniques to address the unique challenges inherent in unbounded street perspectives, demonstrating promise for a range of subsequent downstream tasks. While advancements in ray marching and enhanced MLPs have been notable, these methods remain a trade-off between rendering fidelity and efficiency, which may restrict their utility in expansive autonomous driving scenarios.
@@
3DGS (Kerbl et al. 2023) is a seminal work that pioneers the use of 3D Gaussian ellipsoids to explicitly represent a scene and leverages CUDA for parallel rendering, thereby surpassing NeRFs in both rendering quality and efficiency. HUGS, as introduced in (Zhou et al. 2024a), enables real-time rendering of new perspectives, delivering both 2D and 3D semantic data, delivering both 2D and 3D semantic information with remarkable precision.
@@
Figure 2: The pipeline of the proposed method for driving scene reconstruction. Given consecutive multi-camera images along with their respective road and non-road masks, we initially generate decoupled road pcd (point cloud) and environment pcd, a road SDF is then pre-trained as subsequent guidance for the road Gaussian model. The Environment pcd enables the initialization for the environment Gaussian model which composes rendered images with images by paratactic road model via the proposed depth-ordered hybrid rendering.
@@
The vanilla 3DGS focuses solely on pixel color consistency, often overlooking the modeling of geometric structures. Consequently, obvious discontinuities such as artifacts and holes appear under novel viewpoints. This discontinuity phenomenon would be more prominent on AD datasets with sparse training views. DRGS (Chung, Oh, and Lee 2024) introduces a dense depth map as geometry guidance to mitigate overfitting. NeuSG (Chen, Li, and Lee 2023) flattens the Gaussian ellipsoid by designing a regularization loss for the scale parameter, and performs consistent joint optimization in combination with the signed distance field. Scaffold GS (Lu et al. 2024) precisely controls the density of the Gaussian ellipsoid by establishing voxel anchors and corresponding offset vectors and learning the various parameters of the Gaussian through a simple multi-layer network. Gaussian Pro (Cheng et al. 2024) considers the planar prior in the scene, explicitly constraining the growth of Gaussians, achieving amendatory rendering and more compact representation. 2DGS (Huang et al. 2024a) presents a highly efficient differentiable 2D Gaussian renderer, enabling perspective-accurate splatting by leveraging 2D surface modeling, demonstrating superiority in geometric reconstruction. Driving Gaussians (Zhou et al. 2024b), Street Gaussians (Yan et al. 2024) and S3 Gaussians (Huang et al. 2024b) all initialize the point cloud collected by LiDAR instead of the point cloud generated by SfM. Leveraging the dense, accurate data from LiDAR, additional supervision, such as the positional constraint, is introduced to the learning process.
As far as we know, Gaussian Splatting works for AD in the literature pay little attention to the capability of novel view synthesis, resulting in low-quality images under significant camera transformations. We aim to address this issue by decoupling the entire scene into road and non-road regions, optimizing each with guidance from hybrid rendering and geometric regularization of Gaussians, achieving more robust and higher-fidelity rendering quality.
@@
where $\mu$ is the mean position of Gaussian and $x$ is a discretional coordinate within the Gaussian space. $\mathbf{\Sigma}=RSS^{T}R^{T}$, where $R$ is the rotation matrix and $S$ is the scaling matrix. Apart from the spherical harmonic coefficients that represent the anisotropic color of each Gaussian, each pixel color is aggregated by $\alpha$-blending:
$\mathbf{c}(\mathbf{x})=\sum_{i=1}^{N}{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
@@
The proposed method utilizes initial points clouds and semantic masks as paratactic input of multi-camera views. The pcd (point cloud) initialization phase first generates road and non-road point clouds, which will be further modeled as road and environment Gaussian model. Based on the known road point cloud, we propose to design a mathematical implicit road representation by SDF which serves as a prior for the surface training. The SDF-based surface constraint consists of pre-training and off-line supervision phases, by taking advantage of the distance and normal characters geometrically. We use distinct Gaussians to model road and non-road elements, enhancing the rendering quality during perspective shifts. To achieve that, depth-ordered hybrid rendering is delicately designed, by which the road surface and non-road areas can be coupled and superimposed consistently and continuously, yielding superior performance compared with SOTA methods that employ individual Gaussian model. The rendered image by fusing two models will be supervised by ground truth via Gaussian loss, and optimized in collaboration with the regularization terms. We will sequentially detail this process in the following sections.
Initialization of the Decoupled PCD
@@
Distinguishing ourselves from prior methods, we opt to utilize LiDAR scan which exhibits better multi-view consistency and prior geometric structure compared to the point cloud from SfM. The initial point cloud is classified into road and non-road parts for the subsequent training. As shown in Figure 2, based on the calibrated intrinsic and extrinsic parameters, multi-camera images and their masks generated by Mask2Former (Cheng et al. 2022) are utilized to form colored and semantically labeled single-frame point clouds. Then we stitch these separate point clouds to build road and environment point clouds.
Pre-trained Surface Based On SDF
@@
To optimize the geometric structure of roads and maintain continuity when facing obvious changes in perspective, we design a road constraint based on surface guidance. Utilizing the segregated road point cloud, we pre-train a signed distance field to represent the road surface implicitly. This approach differs from the joint ray sampling optimization in NeRF and the guidance of geometric structure and image consistency in GSDF (Yu et al. 2024). As road point clouds collected by LiDAR have been calibrated precisely in advance, the excellent geometric structure is suitable to be prior for road learning. A straightforward strategy is directly applying distance constraints to Gaussians that belong to the road. Whereas, considering the spatial geometric characteristics of ellipsoids, imposing constraints only on the centers of Gaussian ellipsoids cannot fully restore their correct geometric shapes, potentially restricting the Gaussian distribution to some extent. We therefore adopt a pre-trained model to optimize the Gaussian distribution by adjusting Gaussian parameters nearing the road surface. To this end, we design specific regularization terms for road-related Gaussians. These include distance and normal direction constraints that ensure Gaussians stay close to the road surface and align with the normal direction of the road surface concurrently.
@@
We utilize a predefined network $f_{\theta}(x)$ for predicting the corresponding SDF value given the positional coordinate of the point cloud. Data normalization is imposed on all the training and testing coordinates to ensure stability and enhanced performance. To equalize the number of points on and off the isosurface, we define the road point cloud as $\mathcal{P}_{r}$, and assign a hypothetical SDF value of zero to each point $x_{p}\in\mathcal{P}_{r}$ on the isosurface. For points off the isosurface, we establish a proximity range around each point in $\mathcal{P}_{r}$, perform random sampling within this range to obtain sample points $\mathcal{P}_{s}\supset\mathcal{P}_{r}$, for $\forall x_{s}\in\mathcal{P}_{s}$, we build a map from point $x_{s}$ to the signed euclidean distance $d_{s}$ as:
@@
where $x_{r}^{s}$ is the nearest point of $x_{s}$ in $\mathcal{P}_{r}$. Considering that bumps exist in uneven road surfaces, we generate the normal direction for each point in $\mathcal{P}_{r}$ to assist in the training of the SDF. To be specific, for $\forall x_{r}\in\mathcal{P}_{r}$, set $X_{r}$ as $k$-nearest neighbors of $x_{r}$, by performing SVD:
@@
in which the first and last terms refer to SDF value regularization and eikonal regularizer respectively, the middle term represents normal loss, and ||$\cdot$|| ⋅ | refers to the number of points in the point cloud. The optimal parameters $\theta^{*}=\arg\min_{\theta}{L(\theta)}$ will be frozen and used to guide the learning of road Gaussians in the following processes.
@@
Table 1: The quantitative comparison between our method with four SOTA works on PSNR, SSIM, LPIPS, and FID metrics computed over all clips on the Waymo dataset. We adopt the same training data for each method and follow the corresponding experimental settings strictly. The best and second best are marked in bold and underlined for each metric respectively.
Figure 3: The diagram illustrates the proposed depth-ordered hybrid rendering strategy for the environment and road model. Corresponding primitives of each model generate pixels with colors independently through Gaussian splatting. These colors are then composited based on their rendered depths and transmittances, producing the final rendered image.
@@
where $\lambda_{r}$ and $\lambda_{e}$ are the weights for fusing $I_{r}$ and $I_{e}$ rendered from the environment model and road model respectively, which are organized as:
@@
Figure 4: Comparison of different methods on the Waymo dataset, the left column and right column display the quality of scene reconstruction and novel view synthesis respectively. Our method achieves high-quality reconstruction for both the environment region and the road areas, excelling over other comparative methods in these aspects.
@@
where ${M}$ refers to the semantic mask of the image, ||$\cdot$|| ⋅ | denotes the total number of pixels for given matrix. The $\mathcal{L}_{c}$ enforces consistency between the depth rendered by environment Gaussians and road Gaussians:
@@
where $i$ and $j$ satisfying $\tilde{M}^{i,j}\not=0$, $\tilde{M}$ refers to the banded boundary generated by binary mask $M$, and the orange band at the bottom right corner of Figure 2 intuitively shows the distribution of elements in $\tilde{M}$. We follow Plenoxels (Fridovich-Keil et al. 2022) and use a total variation loss $\mathcal{L}_{\text{tv}}$. The collective impact of the aforementioned loss functions enables DHGS to generate high-quality renderings that preserve a continuous geometric structure.
@@
Datasets and Metrics. We experiment on the publicly available urban scene datasets Waymo (Sun et al. 2020), 32 static scenes are adopted as the same as EmerNeRF (Yang et al. 2023a). We employ multiple metrics to evaluate the performance of the proposed method and baselines. PSNR, SSIM, and LPIPS metrics are utilized to validate the quality of rendered images with ground truth. For novel views without ground truth, we introduce FID (Heusel et al. 2017) to evaluate the image quality. The averaged metrics across all clips are calculated as comparative values.
Baselines. We compare our method with state-of-the-art and publicly available works. 3DGS (Kerbl et al. 2023), 2DGS (Huang et al. 2024a), Gaussian Pro (Cheng et al. 2024), and Scaffold GS (Lu et al. 2024) are chosen as the comparison methods having been regarded for their outstanding performances.
Figure 5: Visual comparisons on the free-view novel view synthesis. The left and right columns respectively exhibit the results of Set3 and Set4 under the free viewpoint setting, where our method significantly outperforms other comparative methods in capturing both road and environment details.
Implementation details. The initial point cloud is randomly downsampled to 600,000 points. All methods are trained for 60,000 iterations. For the pre-trained SDF network, we employ an 8-layer MLP and train it for 20,000 iterations. The normals of surface points are computed using SVD with 50 neighbors, and the weights for the normal loss and the eikonal loss of the SDF network are set to 0.1 and 0.01. The weights $\lambda_{\text{tran}}$, $\lambda_{\text{sdf}}$, $\lambda_{\text{cons}}$, $\lambda_{\text{tv}}$, $\lambda_{d}$ and $\lambda_{n}$ of the overall training loss are set to 0.1, 1.0, 0.04, 0.1, 0.1 and 0.1. We set the densification intervals for the Gaussian ellipsoids in the environment model and road model to 200 and 300 respectively. The learning rate for the anchor position in the environment model is set as 0.00016, the remaining learning rates are kept consistent with the original settings. For each clip, we utilize all the front, left-front, and right-front views of the entire sequence for training, with a 4:1 split ratio between the training and testing sets. The training set is referred to as the reconstruction views, while the testing set is called the novel views. We further apply unrestricted transformations on camera poses to validate the potential of the model on data-driven simulation, which is referred to as free-view novel view synthesis.
@@
We benchmark our method against established baselines in scene reconstruction, novel view synthesis, and free-view novel view synthesis that stands for notable viewpoints transfer. All methods employ a consistent set of reconstruction views for training, while the remainder are employed as novel views for testing in each clip. The qualitative and quantitative results will be examined in the subsequent sections, with an ablation study included for analysis.